# A mesh-free hybrid Chebyshev-Tucker tensor format with applications to multi-particle modelling

### v1

@@ 3.1.2 Algebraic case with full tensor input
In many applications, see [30, 26, 5] among others, including ours on multi-particle modelling which will be discussed in Section 4, the function $f$ may not be given explicitly in the full domain $[-1,1]^{3}$, but only the data on equi-spaced grid are available. Specifically, we only have access to the function values at the uniform grid points collected in a tensor
${\bf F}_{i_{1},i_{2},i_{3}}=f(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)}),\quad i_{\ell}=1,\cdots,n_{\ell},\quad\ell=1,2,3,$ (3.10)
where $t_{i_{\ell}}^{(\ell)}=-1+(i_{\ell}-1)h_{\ell}$ is the grid point and $h_{\ell}=2/(n_{\ell}-1)$ is the grid size.
In this case, to obtain the function values at Chebyshev nodes, we first construct a multivariate cubic spline interpolation $Q(x_{1},x_{2},x_{3})$ of the data ${\bf F}$ on the uniform grid, i.e., $Q(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})={\bf F}_{i_{1},i_{2},i_{3}}$, and then evaluate $Q$ at Chebyshev nodes, as shown in Algorithm 2. The rest of the algorithm remains the same as Algorithm 1. Note that here and forehand, we assume $n_{\ell}\geq 4$ naturally to ensure the existence of the cubic spline.
@@
1:Given a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ discretizing function $f:[-1,1]^{3}\to\mathbb{R}$ on the uniform grid, the Chebyshev degrees ${\bf m}=(m_{1},m_{2},m_{3})$ and tuck_als tolerance $\varepsilon>0$.
@@
4:Evaluate $Q$ at Chebyshev nodes to obtain ${\bf T}$, i.e., ${\bf T}_{i_{1},i_{2},i_{3}}=Q(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)})$.
5:Perform lines 4-7 in Algorithm 1.
@@ 3.1.3 Algebraic case with CP tensor input
In fact, the multivariate cubic spline interpolation, see e.g. [13, Chapter 17], is expensive in terms of computational cost. However, in many applications of interest, e.g., quantum chemistry [26], multi-particle modelling [30] and the discretization of certain differential operators [24], the target input tensor ${\bf F}$ has some inherent low-rank structure and thus can be cheaply represented in some low-rank tensor format. In such cases, the computation of the evaluation tensor ${\bf T}$ can be done more efficiently. In particular, we consider here the case where ${\bf F}$ is given in the CP format, i.e.,
${\bf F}=\sum_{k=1}^{R}\xi_{k}{\bf a}_{k}^{(1)}\otimes{\bf a}_{k}^{(2)}\otimes{\bf a}_{k}^{(3)},$ (3.11)
where ${\bf a}_{k}^{(\ell)}\in\mathbb{R}^{n_{\ell}}$ are $\ell_{2}$ normalized canonical vectors of each mode. In this case, we only need to construct univariate cubic spline interpolations of the canonical vectors ${\bf a}_{k}^{(\ell)}$ in each mode and make tensor products and summations of them. Specifically, let $q_{k}^{(\ell)}(x_{\ell})$ be the cubic spline of the data ${\bf a}_{k}^{(\ell)}$ on the uniform grid $\{t_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{n_{\ell}}$, i.e., $q_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})={\bf a}_{k}^{(\ell)}(i_{\ell})$ and then construct the sum of the tensor product of the univariate splines as
$Q(x_{1},x_{2},x_{3})=\sum_{k=1}^{R}\xi_{k}q_{k}^{(1)}(x_{1})q_{k}^{(2)}(x_{2})q_{k}^{(3)}(x_{3}).$ (3.12)
Finally, the function evaluation tensor ${\bf T}$ can be obtained by evaluating $Q$ at Chebyshev nodes. The rest of the algorithm remains the same as Algorithm 1. This procedure is summarized in Algorithm 3.
@@
1:Given a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ in CP format Eq. 3.11 discretizing function $f:[-1,1]^{3}\to\mathbb{R}$ on the uniform grid, the Chebyshev degrees ${\bf m}=(m_{1},m_{2},m_{3})$ and tuck_als tolerance $\varepsilon>0$.
@@
3:Construct $Q(x_{1},x_{2},x_{3})$ by Eq. 3.12. Evaluate $Q$ at Chebyshev nodes to obtain ${\bf T}$.
4:Perform lines 4-7 in Algorithm 1.
In the case of CP tensor input, one can do better than the above Algorithm 3 by avoiding the construction of the function evaluation tensor ${\bf T}$ then computing the Chebyshev coefficients ${\bf C}$ directly by Eq. 3.7, which requires $\mathcal{O}(m^{3})$ storage where we assume $m=\max\{m_{1},m_{2},m_{3}\}$. In fact, the function evaluations ${\bf T}$ in Algorithm 3 also admits a CP representation given by
${\bf T}_{i_{1},i_{2},i_{3}}=Q(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)})=\sum_{k=1}^{R}\xi_{k}{\bf q}_{k}^{(1)}(i_{1}){\bf q}_{k}^{(2)}(i_{2}){\bf q}_{k}^{(3)}(i_{3})$ (3.13)
where ${\bf q}_{k}^{(\ell)}\in\mathbb{R}^{m_{\ell}}$ contains the function evaluations of the univariate spline $q_{k}^{(\ell)}(x)$ at $\{s_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{m_{\ell}}$, i.e., ${\bf q}_{k}^{(\ell)}(i_{\ell})=q_{k}^{(\ell)}(s_{i_{\ell}}^{(\ell)})$. Therefore, the Chebyshev coefficient tensor ${\bf C}$ can be computed directly by Eq. 3.7 as
${\bf C}={\bf T}\times_{1}W^{(1)}\times_{2}W^{(2)}\times_{3}W^{(3)}=\sum_{k=1}^{R}\xi_{k}{\bf c}_{k}^{(1)}\otimes{\bf c}_{k}^{(2)}\otimes{\bf c}_{k}^{(3)},$ (3.14)
where ${\bf c}_{k}^{(\ell)}=W^{(\ell)}{\bf q}_{k}^{(\ell)}$ for $\ell=1,2,3$. From here, we can leverage the reduced HOSVD (RHOSVD) [31] method instead of full tuck_als to compress ${\bf C}$ directly from canonical to Tucker format, without forming the full tensor ${\bf C}$. The rest of the algorithm remains the same as Algorithm 1. This procedure is summarized in Algorithm 4.
Algorithm 4 ChebTuck Approximation: algebraic case (CP tensor input, RHOSVD)
1:Given a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ in CP format discretizing function $f:[-1,1]^{3}\to\mathbb{R}$ on the uniform grid, the degree of Chebyshev polynomials in each dimension $m_{1},m_{2},m_{3}$ and rhosvd tolerance $\varepsilon>0$.
2:The ChebTuck format $\hat{f}_{\bf m}$ of $f$.
3:Compute the splines $q_{k}^{(\ell)}(x_{\ell})$ and the evaluations ${\bf q}_{k}^{(\ell)}$ for $k=1,\cdots,R$ and $\ell=1,2,3$.
4:Compute the canonical vectors ${\bf c}_{k}^{(\ell)}=W^{(\ell)}{\bf q}_{k}^{(\ell)}$ of ${\bf C}$ for $k=1,\cdots,R$ and $\ell=1,2,3$.
5:Compute the reduced Tucker decomposition: $[\bm{\beta},V^{(1)},V^{(2)},V^{(3)}]=\texttt{rhosvd}({\bf C},\varepsilon)$
6:Perform lines 6-7 in Algorithm 1.
Remark 3.2.
(Mesh-free method) Our Algorithms outline the computational schemes for low-rank tensor approximation of the function $f$ in both the functional and algebraic (discrete) cases. We underline that in both cases our Algorithms are implemented as mesh-free computational schemes in the sense that one needs only access to the functional values on the Chebyshev interpolation grid in 3D (in general, approximate values), but does not require storage of the discretized function on the fine 3D-grid and further computing the Tucker approximation of arising large 3D tensor array by using the Tucker-ALS scheme or by other means (as for the traditional grid based tensor approximation methods). Instead, we suggest to compute the only Tucker-ALS approximation of rather small Chebyshev coefficients tensor (CCT) whose size has no relations with the grid size, required for accurate discretization of the function $f$. In case of CP canonical input we perform only RHOSVD approximation of the CCT instead of Tucker-ALS of full format CCT.
Remark 3.3.
We note here that the Chebyshev coefficient tensor ${\bf C}$ in Algorithm 3 that is the input for tuck_als is mathematically equal to the tensor ${\bf C}$ in Algorithm 4 treated by algorithm RHOSVD while the latter leverages the CP structure of ${\bf C}$ and the memory and computational cost are both $\mathcal{O}(mR)$. Therefore, Algorithm 4 is always preferred in case of CP tensor input. See also Section 4.2.1 for the numerical comparison of the two algorithms.
@@ 3.2.1 Functional case
We first consider the ChebTuck approximation $\hat{f}_{\bf m}$ of a function $f$ obtained by Algorithm 1, i.e. the functional case. It is well know that multivariate Chebyshev interpolation $\tilde{f}_{\bf m}$ converges exponentially fast to the function $f$ in the uniform norm over $[-1,1]^{3}$ with respect to the degree of the Chebyshev polynomials ${\bf m}=(m_{1},m_{2},m_{3})$, provided that $f$ has sufficient regularity. Our ChebTuck format introduces an additional error due to the Tucker approximation of the Chebyshev coefficient tensor ${\bf C}$, which is controlled by the Tucker-ALS truncation error $\varepsilon$. In the following proposition, we show that the ChebTuck approximation $\hat{f}_{\bf m}$ is almost as accurate as the full Chebyshev interpolation $\tilde{f}_{\bf m}$ provided that the Tucker approximation error $\varepsilon$ is sufficiently small.
@@
$\|f-\hat{f}_{\bf m}\|_{\infty}\leq\|f-\tilde{f}_{\bf m}\|_{\infty}+\sqrt{m_{1}m_{2}m_{3}}\varepsilon\|\bf{C}\|_{\mathrm{F}},$ (3.15)
where $\|\cdot\|_{\infty}$ is the uniform norm over $[-1,1]^{3}$.
Proof.
By triangle inequality, it suffices to prove
$\|\hat{f}_{\bf m}-\tilde{f}_{\bf m}\|_{\infty}\leq\sqrt{m_{1}m_{2}m_{3}}\varepsilon\|\bf{C}\|_{\mathrm{F}}.$ (3.16)
Indeed, by the definition of the ChebTuck format Eq. 3.9 and the fact that Chebyshev polynomials are bounded by 1, we have
$\|\tilde{f}_{\bf m}-\hat{f}_{\bf m}\|_{\infty}$ $=\left\|\sum_{i_{1}=1}^{m_{1}}\sum_{i_{2}=1}^{m_{2}}\sum_{i_{3}=1}^{m_{3}}({\bf C}_{i_{1},i_{2},i_{3}}-\hat{\bf C}_{i_{1},i_{2},i_{3}})T_{i_{1}-1}(x_{1})T_{i_{2}-1}(x_{2})T_{i_{3}-1}(x_{3})\right\|_{\infty}$ (3.17)
$\leq\sum_{i_{1}=1}^{m_{1}}\sum_{i_{2}=1}^{m_{2}}\sum_{i_{3}=1}^{m_{3}}|{\bf C}_{i_{1},i_{2},i_{3}}-\hat{\bf C}_{i_{1},i_{2},i_{3}}|$
$\leq\sqrt{m_{1}m_{2}m_{3}}\|{\bf C}-\hat{\bf{C}}\|_{\mathrm{F}}\leq\sqrt{m_{1}m_{2}m_{3}}\varepsilon\|{\bf C}\|_{\mathrm{F}},$
where we also used the Cauchy-Schwarz inequality and the relative truncation error $\varepsilon$ of the Tucker approximation in the last line. ∎
@@ 3.2.2 Algebraic case
Now, we consider the ChebTuck approximation in the algebraic case, i.e., given a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ discretizing a black-box function $f$ on the uniform grid. In this case, the metric for the error of the approximation is the difference between the ChebTuck approximation $\hat{f}_{\bf m}$ evaluated at the uniform grid and the tensor ${\bf F}$
$\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|.$ (3.18)
In other words, we hope to recover the original input tensor ${\bf F}$ when we evaluate the ChebTuck format $\hat{f}_{\bf m}$ at the uniform grid. Since we know nothing about the function $f$ except for the values on the grid, for Algorithm 2, we can say nothing about the error of the Chebyshev interpolation of the multivariate cubic spline $Q$, thus neither the ChebTuck approximation $\hat{f}_{\bf m}$. For example, imagine that ${\bf F}$ is a tensor of random numbers, then we would not expect the Chebyshev interpolation thus also the ChebTuck approximation to be accurate.
Fortunately, in the case of CP tensor input, we can provide a bound for the error of the ChebTuck approximation $\hat{f}_{\bf m}$ depending on the errors of the univariate Chebyshev interpolations of the canonical vectors ${\bf a}_{k}^{(\ell)}$ in each mode. Note that the canonical vectors ${\bf c}_{k}^{(\ell)}$ of the Chebyshev coefficient tensor ${\bf C}$ are nothing but the univariate Chebyshev coefficients of the splines $q_{k}^{(\ell)}(x)$. In other words, let $g_{k}^{(\ell)}(x):=\sum_{j=1}^{m_{\ell}}{\bf c}_{k}^{(\ell)}(j)T_{j-1}(x)$, then $g_{k}^{(\ell)}(x)$ is the Chebyshev interpolation of $q_{k}^{(\ell)}(x)$. Let $\hat{\bf C}$ be the RHOSVD truncation of ${\bf C}$, then the ChebTuck approximation error is given by
$\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|+\|\hat{f}_{\bf m}-\tilde{f}_{\bf m}\|_{\infty}$ (3.19)
where $\tilde{f}_{\bf m}$ is the Chebyshev interpolant Eq. 3.1 with coefficients ${\bf C}$ given in Eq. 3.14, i.e., before RHOSVD truncation. $\tilde{f}_{\bf m}$ is thus given by
$\tilde{f}_{\bf m}(x_{1},x_{2},x_{3})=\sum_{k=1}^{R}\xi_{k}g_{k}^{(1)}(x_{1})g_{k}^{(2)}(x_{2})g_{k}^{(3)}(x_{3}).$ (3.20)
Then, we have the following error bound for the first term in the right-hand side of Eq. 3.19.
@@
Let $\tilde{f}_{\bf m}$ be of the form Eq. 3.1 with coefficients ${\bf C}$ given by Algorithm 4 before RHOSVD truncation, i.e., $\tilde{f}_{\bf m}$ takes the form Eq. 3.20. Then we have
$\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\sum_{k=1}^{R}\xi_{k}\sum_{\ell=1}^{3}\prod_{r=1}^{\ell-1}\left(1+\delta_{k}^{(r)}\right)\delta_{k}^{(\ell)},$ (3.21)
where $\delta_{k}^{(\ell)}:=\|q_{k}^{(\ell)}-g_{k}^{(\ell)}\|_{\infty}$ is the error of the univariate Chebyshev interpolation of $q_{k}^{(\ell)}(x)$. In particular, let $\delta=\max_{k,\ell}\delta_{k}^{(\ell)}$, then
$\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\|\bm{\xi}\|_{1}\left((1+\delta)^{3}-1\right),$ (3.22)

### v2

@@ 3.1.2 Algebraic case with full tensor input
In many applications, see [37, 34, 7] among others, including ours on multi-particle modeling which will be discussed in Section 4, the function $f$ may not be given explicitly in the full domain $1,1]^{3}$, but only the data on the equi-spaced grid are available. Specifically, we only have access to the function values at uniform grid points collected in a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, ${\bf F}_{i_{1},i_{2},i_{3}}=f(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})$ for $i_{\ell}=1,\ldots,n_{\ell}$, $\ell=1,2,3$. Here $t_{i_{\ell}}^{(\ell)}=-1+(i_{\ell}-1)h_{\ell}$ is the grid point and $h_{\ell}=2/(n_{\ell}-1)$ is the grid size.
In this case, we first extrapolate the function to the full domain $1,1]^{3}$ by constructing a multivariate cubic spline interpolation $Q(x_{1},x_{2},x_{3})$ of the data $\bf F$} on the uniform grid, i.e., $Q(x_{1},x_{2},x_{3})$ is a $C^{2}$ piecewise cubic polynomial satisfying the interpolatory property $Q(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})={\bf F}_{i_{1},i_{2},i_{3}}$, and then call Algorithm 1 with $f=Q$. This procedure is summarized in Algorithm 2. Note that here and forehand, we assume $n_{\ell}\geq 4$ naturally to ensure the existence of the cubic spline.
@@
1:Given a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathbf{m}$, $\varepsilon>0$ as in Algorithm 1.
@@
4:Call Algorithm 1 with $f=Q$.
@@ 3.1.3 Algebraic case with CP tensor input
In fact, the multivariate cubic spline interpolation, see e.g. [17, Chapter 17], is computationally expensive. However, in many applications of interest, e.g., quantum chemistry [34], multi-particle modeling [37] and the discretization of certain differential operators [32], the target input tensor $\bf F$} has some inherent low-rank structure and thus can be cheaply represented in some low-rank tensor format. In such cases, the computation of the evaluation tensor $\bf T$} can be done more efficiently. In particular, we consider here the case where $\bf F$} is given in the CP format, i.e.,
(18) ${\bf F}={\sum}_{k=1}^{R}\xi_{k}{\bf a}_{k}^{(1)}\otimes{\bf a}_{k}^{(2)}\otimes{\bf a}_{k}^{(3)},$
where ${\bf a}_{k}^{(\ell)}\in\mathbb{R}^{n_{\ell}}$ are $\ell_{2}$ normalized canonical vectors of each mode. In this case, we only need to construct univariate cubic splines of the canonical vectors ${\bf a}_{k}^{(\ell)}$ in each mode, and make tensor products and summations of them. Specifically, let $q_{k}^{(\ell)}(x_{\ell})$ be the cubic spline of the data ${\bf a}_{k}^{(\ell)}$ on the uniform grid $\{t_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{n_{\ell}}$, i.e., $q_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})={\bf a}_{k}^{(\ell)}(i_{\ell})$ and then construct the sum of the tensor product of the univariate splines as
(19) $Q(x_{1},x_{2},x_{3})={\sum}_{k=1}^{R}\xi_{k}q_{k}^{(1)}(x_{1})q_{k}^{(2)}(x_{2})q_{k}^{(3)}(x_{3}).$
Calling Algorithm 1 with $f=Q$, one sees immediately from the definitions of the CCT $\bf C$} Eqs. 13 and 15 that $\bf C$} computed in line 3 is also in CP format. Specifically, let ${\bf q}_{k}^{(\ell)}\in\mathbb{R}^{m_{\ell}}$ contain the function evaluations of the univariate spline $q_{k}^{(\ell)}(x)$ at $\{s_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{m_{\ell}}$, i.e., ${\bf q}_{k}^{(\ell)}(i_{\ell})=q_{k}^{(\ell)}(s_{i_{\ell}}^{(\ell)})$, then by definition of the function evaluation tensor $\bf T$} in Eq. 13 and the CCT $\bf C$} in Eq. 15, we have
(20) ${\bf T}={\sum}_{k=1}^{R}\xi_{k}{\bf q}_{k}^{(1)}\otimes{\bf q}_{k}^{(2)}\otimes{\bf q}_{k}^{(3)}\implies{\bf C}={\sum}_{k=1}^{R}\xi_{k}{\bf c}_{k}^{(1)}\otimes{\bf c}_{k}^{(2)}\otimes{\bf c}_{k}^{(3)},$
with ${\bf c}_{k}^{(\ell)}=W^{(\ell)}{\bf q}_{k}^{(\ell)}$ for $\ell=1,2,3$ and $W^{(\ell)}$ being the inverse DCT matrix defined in Eq. 14. From here, we can leverage the reduced HOSVD (rhosvd) [39] method instead of full tuck_als to compress $\bf C$} directly from canonical to Tucker format, without forming the full tensor $\bf C$}. The rest of the algorithm remains the same as Algorithm 1. This procedure is summarized in Algorithm 3.
@@
1:Given a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ in CP format, the Chebyshev degrees ${\bf m}=(m_{1},m_{2},m_{3})$ and rhosvd tolerance $\varepsilon>0$.
@@
3:Compute the splines $q_{k}^{(\ell)}(x_{\ell})$ and the evaluations ${\bf q}_{k}^{(\ell)}$ for $k=1,\ldots,R$, $\ell=1,2,3$.
4:Compute the canonical vectors ${\bf c}_{k}^{(\ell)}=W^{(\ell)}{\bf q}_{k}^{(\ell)}$ of $\bf C$} for $k=1,\ldots,R$, $\ell=1,2,3$.
5:Compute the reduced Tucker decomposition: $\bm{\beta},V^{(1)},V^{(2)},V^{(3)}]=\texttt{rhosvd}({\bf C},\varepsilon)$
6:Perform lines 5-6 in Algorithm 1.
@@ 3.2 Error bounds for the Chebyshev-Tucker approximation
In this section, we always assume $m=m_{1}=m_{2}=m_{3}$ without loss of generality to simplify the notation. While we state the results for general tensor order $d$ since they also hold for $d\geq 3$, the proofs specialize to $d=3$ for clarity of presentation.
@@ 3.2.1 Functional case
We first consider the ChebTuck approximation $\hat{f}_{\bf m}$ of a function $f$ obtained by Algorithm 1, i.e. the functional case. It is well-known that multivariate Chebyshev interpolation $\tilde{f}_{\bf m}$ converges exponentially fast to the function $f$ in the uniform norm over $1,1]^{3}$ with respect to the degree $\bf m$} of the Chebyshev polynomials, provided that $f$ has sufficient regularity. Our ChebTuck format introduces an additional error due to the Tucker approximation of the Chebyshev coefficient tensor $\bf C$}, which is controlled by the Tucker-ALS truncation error $\varepsilon$. In the following proposition, we show that the ChebTuck approximation $\hat{f}_{\bf m}$ is almost as accurate as the full Chebyshev interpolation $\tilde{f}_{\bf m}$ provided that the Tucker approximation error $\varepsilon$ is sufficiently small.
@@
(21) $\|f-\hat{f}_{\bf m}\|_{\infty}\leq\|f-\tilde{f}_{\bf m}\|_{\infty}+m^{d/2}\varepsilon\|\bf{C}\|_{\mathrm{F}},$
where $\|\cdot\|_{\infty}$ is the uniform norm over $1,1]^{3}$.
Proof 3.3.
By triangle inequality, it suffices to prove $\|\hat{f}_{\bf m}-\tilde{f}_{\bf m}\|_{\infty}\leq m^{3/2}\varepsilon\|\bf{C}\|_{\mathrm{F}}$. Indeed, by the definition of the ChebTuck format Eq. 17 and the fact that Chebyshev polynomials are bounded by 1, we have
$\|\tilde{f}_{\bf m}-\hat{f}_{\bf m}\|_{\infty}$ $=\left\|\sum_{i_{1}=1}^{m}\sum_{i_{2}=1}^{m}\sum_{i_{3}=1}^{m}({\bf C}_{i_{1},i_{2},i_{3}}-\hat{\bf C}_{i_{1},i_{2},i_{3}})T_{i_{1}}(x_{1})T_{i_{2}}(x_{2})T_{i_{3}}(x_{3})\right\|_{\infty}$
$\leq\sum_{i_{1}=1}^{m}\sum_{i_{2}=1}^{m}\sum_{i_{3}=1}^{m}|{\bf C}_{i_{1},i_{2},i_{3}}-\hat{\bf C}_{i_{1},i_{2},i_{3}}|\leq m^{3/2}\|{\bf C}-\hat{\bf{C}}\|_{\mathrm{F}}\leq m^{3/2}\varepsilon\|{\bf C}\|_{\mathrm{F}},$
where we used the Cauchy-Schwarz inequality and the relative truncation error $\varepsilon$ of the Tucker-ALS approximation in the last two inequalities.
@@ 3.2.2 Algebraic case
Now, we consider the ChebTuck approximation in the algebraic case, i.e., given a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ assumed to discretize a black-box function $f$ on the uniform grid. In this case, the metric for the error of the approximation is the difference between the ChebTuck approximation $\hat{f}_{\bf m}$ evaluated at the uniform grid and the tensor $\bf F$}, i.e., $\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|$. In other words, we hope to recover the original input $\bf F$} when we evaluate the ChebTuck format $\hat{f}_{\bf m}$ at the uniform grid. Since we know nothing about the function $f$ except for the values on the grid, for Algorithm 2, we can say nothing about the error of the Chebyshev interpolation of the multivariate cubic spline $Q$, thus neither the ChebTuck approximation $\hat{f}_{\bf m}$. For example, for highly non-regular function $f$ we would not expect the Chebyshev interpolation thus also the ChebTuck approximation to be accurate.
Fortunately, in the case of CP tensor input, we can provide a bound for the error of the ChebTuck approximation $\hat{f}_{\bf m}$ depending on the errors of the univariate Chebyshev interpolations of the canonical vectors ${\bf a}_{k}^{(\ell)}$ in each mode. Note that the canonical vectors ${\bf c}_{k}^{(\ell)}$ of the Chebyshev coefficient tensor $\bf C$} are nothing but the univariate Chebyshev coefficients of the splines $q_{k}^{(\ell)}(x)$. In other words, let $g_{k}^{(\ell)}(x):=\sum_{j=1}^{m}{\bf c}_{k}^{(\ell)}(j)T_{j}(x)$, then $g_{k}^{(\ell)}(x)$ is the Chebyshev interpolation of $q_{k}^{(\ell)}(x)$. Let $\hat{\bf C}$ be the RHOSVD truncation of $\bf C$}, then the ChebTuck approximation error is given by
(22) $\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|+\|\hat{f}_{\bf m}-\tilde{f}_{\bf m}\|_{\infty}$
where $\tilde{f}_{\bf m}$ is the Chebyshev interpolant Eq. 12 with coefficients $\bf C$} given in Eq. 20, i.e., before RHOSVD truncation. $\tilde{f}_{\bf m}$ is thus given as
(23) $\tilde{f}_{\bf m}(x_{1},x_{2},x_{3})={\sum}_{k=1}^{R}\xi_{k}g_{k}^{(1)}(x_{1})g_{k}^{(2)}(x_{2})g_{k}^{(3)}(x_{3}).$
Then we have the following bound for the first term in the right-hand side of Eq. 22.
@@
Let $\tilde{f}_{\bf m}$ be of the form Eq. 12 with coefficients $\bf C$} in Eq. 20, i.e., $\tilde{f}_{\bf m}$ takes the form Eq. 23. Let $\delta_{k}^{(\ell)}:=\|q_{k}^{(\ell)}-g_{k}^{(\ell)}\|_{\infty}$ be the error of the univariate Chebyshev interpolation of $q_{k}^{(\ell)}(x)$ and $\delta:=\max_{k,\ell}\delta_{k}^{(\ell)}$, then we have
(24) $\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(d)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\|\bm{\xi}\|_{1}\left((1+\delta)^{d}-1\right),$